there are 1000 students, 40 teachers, and 5 administrators. if the school grows to 1200 students and the ratios maintained, how many administrators are needed?
Answer by TimothyLamb(4379)

5/1000 = x/1200
x = 1200*5/1000
x = 6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php